A pharmacist wishes to mix a solution that is 8% minoxidil. She has 70 ml of 4% solution and wishes to add some 10% solution to obtain the desired 8% solution. How much 10% should she add?
Answer by TimothyLamb(4379)

x = ml of 4% = 70
y = ml of 10%
---
4x + 10y = 8(x + y)
4x + 10y = 8x + 8y
2y = 4x
y = 2x
y = 2*70
y = 140
---
answer:
add 140 ml of 10%
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php